OpenStudy (anonymous):

Given 2 real numbers whose sum is 14. find the minimum possible value for the sum of their squares (answer is an integer or a reduced fraction; NO GUESSING)

OpenStudy (anonymous):

x+y=14 x^2+y^2= __ 2nd equation is the one we want to minimize substituting first equation into 2nd equation x^2+ (14-x)^2 = ___ take the 2nd equation and find the first derivative and set it equal to zero then solve for x then solve for y then plug back into 2nd original equation to determine lowest value

OpenStudy (anonymous):

Can you be more descriptive with the last part, starting from setting it to zero

OpenStudy (anonymous):

 assume that the function for the 2nd equation looks something like that

OpenStudy (anonymous):

we want to find the lowest point this can be done by taking the first derivative of the of the 2nd equation and setting it equal to zero the first derivative gives you an equation for the slope of the line at any point of the graph so at the lowest point of the graph, the slope is equal to zero so taking the first derivative of the 2nd equation with respect to x and setting it equal to zero gives us a solveable equation

OpenStudy (anonymous):

so when you take first derivative and set it equal to zero you will get something that looks like 2x-28+2x =0 so solve for x and that will give you the x value at the lowest value

OpenStudy (anonymous):

Refer to the attached Mathematica solution: